OpenStudy (elleblythe):

How to solve for Arcsin(cos 5pi/8)???

OpenStudy (anonymous):

\[\sin^{-1}(\cos(\frac{5 \pi}{8}))\]

OpenStudy (anonymous):

You need to convert the inner cos into sin, and you are done.. :)

OpenStudy (aum):

\[ \cos(x) = \sin(\frac \pi 2 - x) \]

OpenStudy (anonymous):

Yeah we can do this also.. :)

OpenStudy (anonymous):

arcsin=-5pi

OpenStudy (anonymous):

cos=5pi/8 usetriangle a2+b2=c2 solver for opp arcsin=opp/adj simplify to get arcsin=-5pi

OpenStudy (aum):

\[ \sin^{-1}(\cos(\frac{5 \pi}{8})) = \sin^{-1}(\sin(\frac \pi 2 - \frac{5 \pi}{8})) = \sin^{-1}(\sin(\frac 4\pi 8 - \frac{5 \pi}{8})) = \\ \sin^{-1}(\sin(-\frac {\pi}{8})) = -\frac{\pi}{8} \]

OpenStudy (anonymous):

If you do the other way pi/2=5pi/8=pi/8, so you get two answers

OpenStudy (anonymous):

negative because arc -pi/8

OpenStudy (anonymous):

so two solutions depend on the formula you use, so -5pi, or -pi/8

OpenStudy (elleblythe):

I just checked and -pi/8 was also the final answer in the problem set that I'm answering but thanks to everyone for your efforts

OpenStudy (anonymous):

it just depends what formula you use, you will get different answers

OpenStudy (elleblythe):

@aum I have a question why did you subtract 5pi/8 from pi/2? is it because of the range of arcsin?

OpenStudy (aum):

-5pi is not the correct answer. sin(-5pi) is not equal to cos(5pi/8)

OpenStudy (aum):

because cos(x) can be written as sin(pi/2 - x)

OpenStudy (anonymous):

formula sin(pi/2-5pi/8) use your fx and solver

OpenStudy (anonymous):

what I did was use a right triangle we know cos adj/opp so solve for opp =8-5pi/8

OpenStudy (aum):

Noticed a typo in an earlier reply: \[ \sin^{-1}(\cos(\frac{5 \pi}{8})) = \sin^{-1}(\sin(\frac \pi 2 - \frac{5 \pi}{8})) = \sin^{-1}(\sin(\frac {4\pi}{ 8} - \frac{5 \pi}{8})) = \\ \sin^{-1}(\sin(-\frac {\pi}{8})) = -\frac{\pi}{8} \]

OpenStudy (anonymous):

sin opp/hyp 8-5pi/8 simplify get 5pi

OpenStudy (anonymous):

aum is using the trig identites, so what ever way is easier for you, do it

OpenStudy (anonymous):

he is using the cofunction identites

OpenStudy (aum):

5pi/8 is an ANGLE. It is not the sides. So the adjacent is not 5pi and the hypotenuse is not 8. 5pi/8 = 112.5 degrees.

OpenStudy (anonymous):

Yeah you are right, if you had sides then you could have done it that way